A person has 20 coins, all nickels and dimes, worth one dollar and 40 cents. How many nickels are there?
If n = number of nickels and d = number of dimes, which system of equations represents the problem?
A.)n + d = 20 and n + d = 1.40
B.)n + d = 20 and 5n + 10d = 1.40
C.)n + d = 20 and 5n + 10d = 140

Found 2 solutions by Edwin McCravy, ikleyn:
Answer by Edwin McCravy(20081)


A.)n + d = 20 and n + d = 1.40

That's wrong because if n + d = 20, it can't equal 1.40 too!  


B.)n + d = 20 and 5n + 10d = 1.40

That's wrong because the left side of 5n + 10d = 1.40 is in
cents but the right side is in dollars.


C.)n + d = 20 and 5n + 10d = 140

That's right, for the left side of 5n + 10d = 140 is in cents
and the right side is also in cents, because $1.40 equals 140 cents.

Edwin

Answer by ikleyn(53763)

.
Do you know the children story about a wooden boy Chipollino (the other name Buratino), the boy made of wood?

This problem recalls me the chapter of this story, where the baby doll Malvina teaches small Buratino to read and to make simplest arithmetic.


It is good for a children tale story, but is inappropriate for this forum.

We do not teach little wooden boys in this forum. Really not.